# Electromagnetic field in a cavity induced by gravitational waves

### v1

@@ Abstract
The detection method of gravitational waves (GW) using electromagnetic (EM) cavities has garnered significant attention in recent years. This paper thoroughly examines the analysis for the perturbation of the EM field and raises some issues in the existing literature. Our work demonstrates that the rigid condition imposed on the material, as provided in the literature, is inappropriate due to its reliance on a gauge-dependent quantity that cannot be controlled experimentally. Instead, we incorporate elasticity into the material and revise the governing equations for the electric field induced by GWs, expressing them solely in terms of gauge-invariant quantities. Applying these equations to a cylindrical cavity with the TM010 mode, we present the GW antenna pattern for the detector.
@@ I Introduction
Electromagnetic (EM) fields play a crucial role in the observation of gravitational waves (GWs). Interferometer-type GW detectors utilize lasers Abbott et al. (2023), and pulsar timing arrays employ EM pulses from millisecond pulsars to detect GWs The EPTA and InPTA collaborations et al. (2023); Reardon et al. (2023); Agazie et al. (2023); Xu et al. (2023). Recently, the idea of using high-sensitivity EM cavities for GW detection has been receiving considerable attention Herman et al. (2021); Aggarwal et al. (2021); Berlin et al. (2022); Domcke et al. (2022). The principle underlying this method involves resonating EM fields when the frequency of GWs closely matches the resonant frequency of the EM cavity. This approach has the advantage of being relatively easy to try because sensitive cavity experiments are already being conducted worldwide to search for axion dark matter Andrew et al. (2023); Bartram et al. (2021); Backes et al. (2021).
The resonant frequency of an EM cavity is determined by the size of the cavity. For instance, assuming a cavity size of approximately $0.1$ meter would yield a resonant frequency in the GHz range. If we consider binary black holes as the source of GWs, planet-mass binary black holes of about $\sim 10^{-5}M_{\odot}$ would be required to produce GWs at GHz frequencies just before merging Maggiore (2007). Such mass black holes are challenging to form from stars Collaboration (2023), hence, if they exist, they are likely to be primordial black holes formed from the density fluctuations in the early universe. Although direct evidence of black holes of this mass has yet to be found, the study Niikura et al. (2019), that suggests the possibility of planet-mass black holes through micro-gravitational lensing events is noteworthy.
@@
Our paper is structured as follows. In Section II.1, we introduce physical laws in curved spacetime necessary for discussing the induced EM field by GWs. Section II.2 covers the basics of the EM cavity. To provide a helpful pedagogical illustration, Section II.3 presents an example of forced oscillation. In Section III.1, we present the covariant perturbation theory, enabling a concise perturbation analysis. Applying this approach, we discuss perturbations of Minkowski spacetime in Section III.2 without choosing any gauge. The perturbations of elasticity are developed in Section III.3, and those of electromagnetism are discussed in Section III.4. For the inside of the vacuum cavity, we provide the equation for the induced electric field by GWs in Section III.5. Section III.6 addresses the inadequacy of the rigid condition given in Berlin et al. (2022); Domcke et al. (2022). Using corrected equations, we present an antenna pattern for the detector in Section IV.
@@ II.1 Physical Laws in Curved Spacetime
where $G$ is the Einstein tensor and $T$ is the stress-energy of matters. The contracted Bianchi identities yields
@@
Let us consider a material with a vacuum cavity, where the motion of the cavity forms a 4-dimensional volume $\mathcal{W}$ in $\mathcal{M}$. Introduce the unit vector field $v$ for a timelike congruence with the normalization condition
@@
such that its values on the material are 4-velocities of the material elements. The spatial metric $\gamma^{\prime}$ and the volume form $\varepsilon^{\prime}$ orthogonal to $v$ are given by
@@
The stress-energy of the material is decomposed into
@@
where $\alpha$ is the temporal-temporal part, $\beta$ is the temporal-spatial part, and $W$ is taken from the spatial-spatial part of $\chi$. For Eqs. 24 and 25, we utilized Eq. 11.
EM fields $F$ are governed by Maxwell’s equations as
@@
where $\Box\equiv\nabla^{a}\nabla_{a}$ is the D’Alembertian as Eq. (6) in Herman et al. (2021). The electric and magnetic fields with respect to the material elements are given by
@@
When the material is a perfect conductor, we can impose the boundary condition on $\partial\mathcal{W}$, which is the 3-dimensional timelike hypersurface between the cavity and the material, given by
@@ II.2 EM Cavity in Minkowski Spacetime
Consider the EM cavity in Minkowski spacetime, where the material and the EM field are weak enough to ignore changes of Minkowski spacetime. In this case, we can set $u$ as the constant vector field of 4-velocity aligning with our laboratory. Inside the cavity, Eq. 29 by contracting with $u$ becomes the homogeneous wave equation for the electric field. When the material is a perfect conductor, we obtain the stationary solution as
@@
The resonant modes of a cylindrical cavity are categorized into transverse magnetic (TM) and transverse electric (TE) modes. In each category, they have indices $(m,n,p)$ where $m=0,1,\cdots$ is the azimuthal, $n=1,2,\cdots$ is the radial, and $p=0,1,\cdots$ is the longitudinal mode number. A resonant mode $\mathfrak{e}_{mnp}\left(\vec{x}\right)$ for TM or TE modes has the parity symmetry as
$\mathfrak{e}_{mnp}\left(r,\pi+\varphi,-z\right)=\left(-1\right)^{m+p}\mathfrak{e}_{mnp}\left(r,\varphi,z\right),$ (38)
where we introduce the cylindrical coordinate system $\left\{r,\varphi,z\right\}$ whose origin is located at the center of the cylinder. For example, the TM010 mode $\mathfrak{e}_{010}\left(\vec{x}\right)$ is given by
$\mathfrak{e}_{010}\left(r,\varphi,z\right)=\frac{\mathfrak{J}_{0}\left(\omega_{010}r\right)}{\mathfrak{J}_{1}\left(\mathfrak{j}_{0,1}\right)}\hat{z},$ (39)
where $\mathfrak{J}_{n}$ is the Bessel function, $\mathfrak{j}_{0,1}$ is the first zero of the zeroth Bessel function, $\omega_{010}\equiv\mathfrak{j}_{0,1}/R$ is the resonance frequency, and $R$ is the radius of the cylinder. Note that the above is normalized by Eq. 37.
@@ III.1 Covariant Perturbation Theory
Let us delve into covariant perturbation theory to facilitate a clear and concise discussion. Consider a foliation $\mathcal{F}$ by a one-parameter family of perturbed spacetime $\mathcal{M}_{\epsilon}$, where $\epsilon$ is the dimensionless perturbation parameter. We set $\mathcal{M}_{0}$ as the background spacetime. To discuss perturbations, we introduce a one-parameter group of diffeomorphisms $\phi_{\epsilon}:\mathcal{M}_{0}\rightarrow\mathcal{M}_{\epsilon}$. Then, the perturbed quantity for a tensor $X$ is given by the pull-back through $\phi_{\epsilon}$ as $\tilde{X}\left(\epsilon\right)\equiv\phi_{-\epsilon}^{*}X$. When $\tilde{X}\left(\epsilon\right)=O\left(\epsilon^{n}\right)$ for a positive integer $n$, it is convenient to introduce a quantity $Y$ such that $\tilde{X}\left(\epsilon\right)=\epsilon^{n}\tilde{Y}\left(\epsilon\right)$, where $\tilde{Y}\left(\epsilon\right)\equiv\phi_{-\epsilon}^{*}Y=O\left(1\right)$. Then, the perturbed quantity is expanded by
@@
where $Y^{\left(0\right)}=\left.Y\right|_{\epsilon=0}$ is the leading-order value, and $\delta Y$ is the linear perturbation. The linear perturbation is provided by the Lie derivative of the quantity, i.e., $\delta Y=\left[\mathcal{L}_{\upsilon}Y\right]_{\epsilon=0}$, where $\upsilon$ is the 5-dimensional vector field on $\mathcal{F}$ generating $\phi_{\epsilon}$. Refer to Figure 1 in Park (2022) for a helpful visualization of the concept.
@@
In the process of determining $\delta Y$ from a measurement, the background spacetime $\mathcal{M}_{0}$ and the leading-order value $Y^{\left(0\right)}$ in $\mathcal{M}_{0}$ are typically known, and the experimental measurement $Y$ is performed in the perturbed spacetime $\mathcal{M}_{\epsilon}$. To determine $\delta Y$ from Eq. 44, ignoring higher-order terms, one needs to choose a gauge to specify the perturbed value $\tilde{Y}\left(\epsilon\right)$. Due to the absence of a preferred gauge, $\delta Y$ cannot be uniquely determined when it is gauge-dependent, making it not measurable. Therefore, gauge invariance for the linear perturbation is essential to enable its measurement. According to the lemma from Stewart and Walker (1974), $\delta Y$ is gauge-invariant if and only if $Y^{\left(0\right)}$ is zero, a constant scalar, or constructed by the Kronecker delta with constant coefficients.
@@
where $\upsilon$ is the generating vector field for $\phi_{\epsilon}$. Utilizing the commutativity of Lie derivative and exterior derivative for the scalar field, we find the perturbation of coordinate dual basis $\left\{dx^{\alpha}\right\}$ as
@@
Introducing pull-backs $\tilde{x}^{\alpha}\left(\epsilon\right)$ and $\tilde{x}^{\prime\alpha}\left(\epsilon\right)$ for $\left\{x^{\alpha}\right\}$ and $\left\{x^{\prime\alpha}\right\}$, respectively, by $\phi_{\epsilon}$, we get the relation given by
@@ III.2 Perturbation of Minkowski Spacetime
Henceforth, we assume that all quantities exist in $\mathcal{M}_{0}$ unless explicitly stated. We also omit the superscript $\left(0\right)$ for brevity when referring to leading-order quantities. Defining $\mathcal{M}_{0}$ as Minkowski spacetime, we have the flat metric $g$, its associated Levi-Civita connection $\nabla$, and the vanishing Riemann tensor $R=0$ at the leading-order. From Eq. 50, the linear perturbation of the Riemann tensor is then given by
@@
which is gauge-invariant due to its vanishing leading-order. Utilizing the commutativity of perturbation and self-contraction, we derive the perturbation of the Ricci tensor as
@@
Assuming the perturbed stress-energy as $\tilde{T}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$, the perturbation of Einstein’s equations in Eq. 1 is given by
@@
Taking the trace of the above equation reveals the vanishing the perturbation of the Ricci scalar, leading to the conclusion that the perturbation of the Ricci tensor also vanishes.
Applying the D’Alembertian to Eq. 57 and utilizing Einstein’s equations from Eq. 60, we obtain the wave equation for the perturbation of Riemann tensor as
@@
where $\left\{t,\vec{x}\right\}$ denotes the globally inertial coordinate system defined in Section II.2, $\omega$ is the parameter for angular frequency, $\kappa$ is the unit spatial vector for the propagation direction, $\int d^{2}\kappa$ represents integration over all directions, $\tilde{R}$ is the amplitude, and $P\left(t,\vec{x};\omega,\kappa\right)\equiv\omega\left(-t+\kappa\cdot\vec{x}\right)$ is the phase. Subsequently, Eq. 57 has the general solution, composed of the particular and homogeneous solutions:
@@
For the geodesic congruence, its unit vector field and extrinsic curvuatre are expanded as follows:
@@
where the leading-order $u$ is the constant 4-velocity aligned to the laboratory, $\delta u$ is its linear perturbation, the leading-order $K$ vanishes, and $\delta K$ is its linear perturbation. The normalization condition Eq. 3 provides
@@
Notice that $\delta K$ is spatial to $u$ and gauge-invariant because its leading-order vanishes. Contracting indices in Eq. 69, we obtain

### v2

@@ Abstract
The detection method of gravitational waves (GW) using electromagnetic (EM) cavities has garnered significant attention in recent years. This paper thoroughly examines the analysis for the perturbation of the EM field and raises some issues in the existing literature. Our work demonstrates that the rigidity condition imposed on the material, as provided in the literature, is inappropriate due to its reliance on a gauge-dependent quantity that cannot be controlled experimentally. Instead, we incorporate elasticity into the material and revise the governing equations for the electric field induced by GWs, expressing them solely in terms of gauge-invariant quantities. Applying these equations to cylindrical cavities, we present the GW antenna patterns for the detector.
@@ I Introduction
Electromagnetic (EM) fields play a crucial role in the observation of gravitational waves (GWs). Interferometer-type GW detectors utilize lasers Abbott et al. (2023), and pulsar timing arrays employ EM pulses from millisecond pulsars to detect GWs EPTA Collaboration and InPTA Collaboration et al. (2023); Reardon et al. (2023); Agazie et al. (2023); Xu et al. (2023). Recently, the ideas of using high-sensitivity EM cavities for GW detection Herman et al. (2021); Aggarwal et al. (2021); Berlin et al. (2022); Domcke et al. (2022) have been receiving considerable attention with their applications Domcke et al. (2023); Berlin et al. (2023); Navarro et al. (2023). The principle underlying this method involves resonating EM fields when the frequency of GWs closely matches the resonant frequency of the EM cavity. This approach has the advantage of being relatively easy to try because sensitive cavity experiments are already being conducted worldwide to search for axion dark matter Yi et al. (2023); Bartram et al. (2021); Backes et al. (2021).
The resonant frequency of an EM cavity is determined by the size of the cavity. For instance, assuming a cavity size of approximately $0.1$ meter would yield a resonant frequency in the GHz range. If we consider binary black holes as the source of GWs, planet-mass binary black holes of about $\sim 10^{-5}M_{\odot}$ would be required to produce GWs at GHz frequencies just before merging Maggiore (2007). Such mass black holes are challenging to form from stars The LVK Collaboration (2023); hence, if they exist, they are likely to be primordial black holes formed from the density fluctuations in the early universe. Although direct evidence of black holes of this mass has yet to be found, the study Niikura et al. (2019), that suggests the possibility of planet-mass black holes through gravitational microlensing events is noteworthy.
@@
Our paper is structured as follows. In Sec. II.1, we introduce physical laws in curved spacetime necessary for discussing the induced EM field by GWs. Sec. II.2 covers the basics of the EM cavity. To provide a helpful pedagogical illustration, Sec. II.3 presents an example of forced oscillation. In Sec. III.1, we present the covariant perturbation theory, enabling a concise perturbation analysis. Applying this approach, we discuss perturbations of Minkowski spacetime in Sec. III.2 without choosing any gauge. The perturbations of elasticity are developed in Sec. III.3, and those of electromagnetism are discussed in Sec. III.4. For the inside of the vacuum cavity, we provide the equation for the induced electric field by GWs in Sec. III.5. Sec. III.6 addresses the inadequacy of the rigid condition given in Berlin et al. (2022); Domcke et al. (2022). Using corrected equations, we present antenna patterns for the detector in Sec. IV.
@@ II.1 Physical laws in curved spacetime
where $G$ is the Einstein tensor and $T$ is the stress–energy tensor of matters. The contracted Bianchi identities yield
@@
Let us consider a material with a vacuum cavity, where the motion of the cavity forms a four-dimensional volume $\mathcal{W}$ in $\mathcal{M}$. Introduce the unit vector field $v$ for a timelike congruence with the normalization condition
@@
such that its values on the material are four-velocities of the material elements. The spatial metric $\gamma^{\prime}$ and the volume form $\varepsilon^{\prime}$ orthogonal to $v$ are given by
@@
The stress–energy tensor of the material is decomposed into
@@
where $\alpha$ is the temporal-temporal part, $\beta$ is the temporal-spatial part, and $W$ is taken from the spatial-spatial part of $\chi$. For Equations 24 and 25, we utilized Eq. 11.
Electromagnetic fields $F$ are governed by Maxwell’s equations as
@@
where $\Box\equiv\nabla^{a}\nabla_{a}$ is the D’Alembertian as Eq. (6) in Herman et al. (2021). The electric field and magnetic fields with respect to the material elements are given by
@@
When the material is a perfect conductor, we can impose the boundary condition on $\partial\mathcal{W}$, which is the three-dimensional timelike hypersurface between the cavity and the material, given by
@@ II.2 EM cavity in Minkowski spacetime
Consider the EM cavity in Minkowski spacetime, where the material and the EM field are weak enough to ignore changes of Minkowski spacetime. In this case, we can set $u$ as the constant vector field of four-velocity aligning with our laboratory. Inside the cavity, Eq. 29 by contracting with $u$ becomes the homogeneous wave equation for the electric field. When the material is a perfect conductor, we obtain the stationary solution as
@@
The resonant modes of a cylindrical cavity are categorized into TM and TE modes. In each category, they have indices $(m,n,p,s)$ where $m=0,1,\cdots$ is the azimuthal and $n=1,2,\cdots$ is the radial mode number. The longitudinal mode number $p$ can be $p=0,1,\cdots$ for TM modes and $p=1,2,\cdots$ for TE modes. For $m\neq 0$, we have two degenerated modes denoted with $s=+1,-1$, respectively. The resonant frequencies for TM and TE modes, respectively, are given by
$\omega^{\mathrm{TM}}_{mnp}$ $=\sqrt{\left(\mathfrak{j}_{mn}/R\right)^{2}+\left(p\pi/L\right)^{2}},$ (38)
$\omega^{\mathrm{TE}}_{mnp}$ $=\sqrt{\left(\mathfrak{j}^{\prime}_{mn}/R\right)^{2}+\left(p\pi/L\right)^{2}},$ (39)
where $R$ and $L$ are the radius and the length of the cylinder, respectively, $\mathfrak{j}_{mn}$ is the $n$-th zero of the Bessel function $\mathfrak{J}_{m}\left(x\right)$, and $\mathfrak{j}^{\prime}_{mn}$ is the $n$-th zero of the derivative of Bessel function $\mathfrak{J}^{\prime}_{m}\left(x\right)$.
In the cylindrical coordinate $\left\{\rho,\phi,z\right\}$ whose origin is located at the center of the cylinder, the resonant modes for TM and TE, respectively, are given by
$\mathfrak{e}^{\mathrm{TM}}_{mnps}\left(\rho,\phi,z\right)=\frac{2\mathfrak{A}^{\mathrm{TM}}_{mnp}}{\mathfrak{J}^{\prime}_{m}\left(\mathfrak{j}_{mn}\right)}\left[\mathfrak{R}^{\mathrm{TM}}_{mn,z}\left(\rho\right)\mathfrak{C}_{-s}\left(m\phi\right)\mathfrak{Z}_{p+}\left(z\right)\hat{z}\right.$
$\qquad\left.-\frac{R}{L}\frac{p\pi}{\mathfrak{j}_{mn}}\left\{\mathfrak{R}^{\mathrm{TM}}_{mn,\rho}\left(\rho\right)\mathfrak{C}_{-s}\left(m\phi\right)\hat{\rho}\left(\phi\right)\right.\right.$
$\qquad\qquad\left.\left.+s\mathfrak{R}^{\mathrm{TM}}_{mn,\phi}\left(\rho\right)\mathfrak{C}_{s}\left(m\phi\right)\hat{\phi}\left(\phi\right)\right\}\mathfrak{Z}_{p-}\left(z\right)\right],$ (40)
$\mathfrak{e}^{\mathrm{TE}}_{mnps}\left(\rho,\phi,z\right)=\frac{2\mathfrak{A}^{\mathrm{TE}}_{mnp}}{\mathfrak{J}_{m}\left(\mathfrak{j}^{\prime}_{mn}\right)}\left[s\mathfrak{R}^{\mathrm{TE}}_{mn,\rho}\left(\rho\right)\mathfrak{C}_{s}\left(m\phi\right)\hat{\rho}\left(\phi\right)\right.$
$\qquad\left.-\mathfrak{R}^{\mathrm{TE}}_{mn,\phi}\left(\rho\right)\mathfrak{C}_{-s}\left(m\phi\right)\hat{\phi}\left(\phi\right)\right]\mathfrak{Z}_{p-}\left(z\right),$ (41)
where $\left\{\hat{\rho}\left(\phi\right),\hat{\phi}\left(\phi\right),\hat{z}\right\}$ is the orthonormal basis of cylindrical coordinate and
$\mathfrak{A}^{\mathrm{TM}}_{mnp}$ $\equiv\left[\left(1+\delta_{m0}\right)\right.$
×$\times\left.\left\{\left(\frac{R}{L}\frac{p\pi}{\mathfrak{j}_{mn}}\right)^{2}\left(1-\delta_{p0}\right)+\left(1+\delta_{p0}\right)\right\}\right]^{-\frac{1}{2}},$ (42)
$\mathfrak{A}^{\mathrm{TE}}_{mnp}$ $\equiv\left[1-\left(m/\mathfrak{j}^{\prime}_{mn}\right)^{2}+\delta_{m0}\right]^{-\frac{1}{2}},$ (43)
$\mathfrak{R}^{\mathrm{TM}}_{mn,\rho}\left(\rho\right)$ $\equiv\mathfrak{J}^{\prime}_{m}\left(\frac{\mathfrak{j}_{mn}}{R}\rho\right),$ (44)
$\mathfrak{R}^{\mathrm{TM}}_{mn,\phi}\left(\rho\right)$ $\equiv\frac{m}{\mathfrak{j}_{mn}\rho/R}\mathfrak{J}_{m}\left(\frac{\mathfrak{j}_{mn}}{R}\rho\right),$ (45)
$\mathfrak{R}^{\mathrm{TM}}_{mn,z}\left(\rho\right)$ $\equiv\mathfrak{J}_{m}\left(\frac{\mathfrak{j}_{mn}}{R}\rho\right),$ (46)
$\mathfrak{R}^{\mathrm{TE}}_{mn,\rho}\left(\rho\right)$ $\equiv\frac{m}{\mathfrak{j}^{\prime}_{mn}\rho/R}\mathfrak{J}_{m}\left(\frac{\mathfrak{j}^{\prime}_{mn}}{R}\rho\right),$ (47)
$\mathfrak{R}^{\mathrm{TE}}_{mn,\phi}\left(\rho\right)$ $\equiv\mathfrak{J}^{\prime}_{m}\left(\frac{\mathfrak{j}^{\prime}_{mn}}{R}\rho\right),$ (48)
$\mathfrak{C}_{s}\left(\phi\right)$$\equiv\begin{cases}\cos\phi&:s=+1\\ \sin\phi&:s=-1\end{cases},$ (49)
$\mathfrak{Z}_{ps}\left(z\right)$ $\equiv\mathfrak{C}_{s}\left(\frac{p\pi}{L}\left(z+\frac{L}{2}\right)\right).$ (50)
For $m=0$, we have to choose $s=-1$ to get a nontrivial mode. Note that the above are normalized by Eq. 37. In addition, these modes have the parity symmetry as
$\mathfrak{e}_{mnps}\left(\rho,\pi+\phi,-z\right)=\left(-1\right)^{m+p}\mathfrak{e}_{mnps}\left(\rho,\phi,z\right).$ (51)
@@ III.1 Covariant perturbation theory
Let us delve into covariant perturbation theory to facilitate a clear and concise discussion. Consider a foliation $\mathcal{F}$ by a one-parameter family of perturbed spacetime $\mathcal{M}_{\epsilon}$, where $\epsilon$ is the dimensionless perturbation parameter. We set $\mathcal{M}_{0}$ as the background spacetime. To discuss perturbations, we introduce a one-parameter group of diffeomorphisms $\phi_{\epsilon}:\mathcal{M}_{0}\rightarrow\mathcal{M}_{\epsilon}$. Then, the perturbed quantity for a tensor $X$ is given by the pullback through $\phi_{\epsilon}$ as $\tilde{X}\left(\epsilon\right)\equiv\phi_{-\epsilon}^{*}X$. When $\tilde{X}\left(\epsilon\right)=O\left(\epsilon^{n}\right)$ for a positive integer $n$, it is convenient to introduce a quantity $Y$ such that $\tilde{X}\left(\epsilon\right)=\epsilon^{n}\tilde{Y}\left(\epsilon\right)$, where $\tilde{Y}\left(\epsilon\right)\equiv\phi_{-\epsilon}^{*}Y=O\left(1\right)$. Then, the perturbed quantity is expanded by
@@
where $Y^{\left(0\right)}=\left.Y\right|_{\epsilon=0}$ is the leading-order value, and $\delta Y$ is the linear perturbation. The linear perturbation is provided by the Lie derivative of the quantity, i.e., $\delta Y=\left[\mathcal{L}_{\upsilon}Y\right]_{\epsilon=0}$, where $\upsilon$ is the five-dimensional vector field on $\mathcal{F}$ generating $\phi_{\epsilon}$. Refer to Fig. 1 in Park (2022) for a helpful visualization of the concept.
@@
In the process of determining $\delta Y$ from a measurement, the background spacetime $\mathcal{M}_{0}$ and the leading-order value $Y^{\left(0\right)}$ in $\mathcal{M}_{0}$ are typically known, and the experimental measurement $Y$ is performed in the perturbed spacetime $\mathcal{M}_{\epsilon}$. To determine $\delta Y$ from Eq. 56, ignoring higher-order terms, one needs to choose a gauge to specify the perturbed value $\tilde{Y}\left(\epsilon\right)$. Owing to the absence of a preferred gauge, $\delta Y$ cannot be uniquely determined when it is gauge dependent, making it not measurable. Therefore, gauge invariance for the linear perturbation is essential to enable its measurement. According to the lemma from Stewart and Walker (1974), $\delta Y$ is gauge invariant if and only if $Y^{\left(0\right)}$ is zero, a constant scalar, or constructed by the Kronecker delta with constant coefficients.
@@
where $\upsilon$ is the generating vector field for $\phi_{\epsilon}$. Utilizing the commutativity of the Lie derivative and exterior derivative for the scalar field, we find the perturbation of coordinate dual basis $\left\{dx^{\alpha}\right\}$ as
@@
Introducing pullbacks $\tilde{x}^{\alpha}\left(\epsilon\right)$ and $\tilde{x}^{\prime\alpha}\left(\epsilon\right)$ for $\left\{x^{\alpha}\right\}$ and $\left\{x^{\prime\alpha}\right\}$, respectively, by $\phi_{\epsilon}$, we get the relation given by
@@ III.2 Perturbation of Minkowski spacetime
Henceforth, we assume that all quantities exist in $\mathcal{M}_{0}$ unless explicitly stated. We also omit the superscript $\left(0\right)$ for brevity when referring to leading-order quantities. Defining $\mathcal{M}_{0}$ as Minkowski spacetime, we have the flat metric $g$, its associated Levi-Civita connection $\nabla$, and the vanishing Riemann tensor $R=0$ at the leading order. From Eq. 62, the linear perturbation of the Riemann tensor is then given by
@@
which is gauge invariant due to its vanishing leading order. Utilizing the commutativity of perturbation and self-contraction, we derive the perturbation of the Ricci tensor as
@@
Assuming the perturbed stress–energy tensor as $\tilde{T}\left(\epsilon\right)=O\left(\epsilon^{2}\right)$, the perturbation of Einstein’s equations in Eq. 1 is given by
@@
Taking the trace of the above equation reveals the vanishing perturbation of the Ricci scalar, leading to the conclusion that the perturbation of the Ricci tensor also vanishes.
Applying the D’Alembertian to Eq. 69 and utilizing Einstein’s equations from Eq. 72, we obtain the wave equation for the perturbation of the Riemann tensor as
@@
where $\left\{t,\vec{x}\right\}$ denotes the globally inertial coordinate system defined in Sec. II.2, $\omega$ is the parameter for angular frequency, $\kappa$ is the unit spatial vector for the propagation direction, $\int d^{2}\kappa$ represents integration over all directions, $\tilde{R}$ is the amplitude, and $P\left(t,\vec{x};\omega,\kappa\right)\equiv\omega\left(-t+\kappa\cdot\vec{x}\right)$ is the phase. Subsequently, Eq. 69 has the general solution, composed of the particular and homogeneous solutions:
@@
For the geodesic congruence, its unit vector field and extrinsic curvature are expanded as follows:
@@
where the leading order $u$ is the constant four-velocity aligned to the laboratory, $\delta u$ is its linear perturbation, the leading order $K$ vanishes, and $\delta K$ is its linear perturbation. The normalization condition Eq. 3 provides
@@
Notice that $\delta K$ is spatial to $u$ and gauge invariant because its leading order vanishes. Contracting indices in Eq. 81, we obtain